OpenStudy (anonymous):

Let f(x) = -4x + 7 and g(x) = 10x - 6. Find f(g(x)).

OpenStudy (anonymous):

My friend, this is mere substitution. If you made g(x) equivalent to u, and you found the value of f at u (which is f(u) = -4u+7), you could then substitute the g(x) back in like this: f(u)=-4u+7=-4g(x)+7=-4(10x-6)+7=-40x+24+7=-40x+31, therefore f(g(x))=-40x+31.